# A Survey on Compositional Learning of AI Models: Theoretical and Experimental Practices

### v1

@@ 3.2 Productivity or Length Generalization
PCFG SET. For the Productivity test [39], the data is split based on sequence lengths. The model is tested on sequences longer than the ones observed during training. For example, in a grammar context, if the model has been trained on the syntax entity-relation-entity, it will be tested on a longer, nested version of this concept, entity-relation-(entity-relation-entity).
CFQ. This is a dataset of Compositional Freebase Questions (CFQ). It is a natural language question-answering task  [49], focusing on semantic parsing, with a SPARQL query against the Freebase knowledge base. There are questions generated at varying levels of complexity. There are various splits available based on input length, output length, input pattern, or output pattern. All these splits aim to maximize compound divergence while minimizing atom divergence. This task seems to test both systematicity and productivity to some extent, though not explicitly, and cannot determine specific areas where a model’s compositional behavior may be lacking.
COGS. From previous explanation, one of the splits of this dataset, Category 3 [Deeper Recursion] [51] is a test for length generalization by increasing the length of the input sequence recursively during testing. Input sequences are generated using nesting of phrases that are longer than those seen during training.
@@ 3.3 Other Generalization Criteria
PCFG SET [39] is the only benchmark that evaluates the other three additional criteria. The Substitutivity or synonymity test [39] uses an input sequence with an atomic unit replaced by a synonymous atomic unit to evaluate how the model prediction changes. Localism is tested by using input sequences composed of smaller sequences A and B. The model is used to translate the full sequences first and then forced to process A and B separately. The outputs of these two experiments are compared to evaluate how local versus global the model is in its compositional reasoning. Overgeneralization test evaluates the model’s results on input sequences that do not conform to the general rules of the dataset, that is, input sequences that are exceptions to the dataset rules. For example, the acquisition of past-tense forms such as the common "ed" ending (open-opened) versus more uncommon forms such as break to broke.
Name Text MM. Sys. Prod. Subst. Loc. Overgen. References PCFG SET  [12] CLUTRR  [27],  [75]  [107] SQOOP  [13], [4] CFQ  [21][33],  [66][8] SCAN  [56],  [81],  [16] COGS  [31][118],  [54] gSCAN  [22],  [111] ReaSCAN  [48] CREPE  [64][106] KiloGram  [55][28] [45] CLEVR  [3], [47],  [80]
@@
4 Compositional Learning Models
Traditional symbolic AI models naturally support compositional reasoning using classical logic applied to formal language understanding [113], formal verification [25], and more. First-order logic can express objects, their properties, and complex compositional relations. Logical operations like conjunction, disjunction, and implication can express compositional structures on which inference rules are applied, supporting complex compositional reasoning [88]. Another classical symbolic formalism includes grammars [10], which can express and generate complex compositional structures. Dealing with noisy and uncertain data is hard with pure symbolic AI. However, probabilistic augmentations and structured output prediction models have been able to explicitly model structural dependencies and support compositional reasoning based on their learned complex patterns from the data [86]. Nevertheless, scalability becomes a challenge for training and inference as the structural dependencies and the number of correlated variables increase. Given these long-lasting challenges of traditional models of compositionality, current neural models have shown success in both scalability and dealing with noisy and sensory data [84]. Especially in modern large language models, complex compositional patterns can be memorized and resemble compositional reasoning. In the rest of this section, we overview the research focused on the development, design, and empirical evaluation of different types of neural models for compositional reasoning.
@@ 4.1 Basic Neural Models
In [39], different neural models are tested on a set of compositional learning tasks. They evaluate Long short-term memory (LSTM) [36], convolutional neural network (CNN) [61], and Transformers for sequence-to-sequence language processing tasks on their proposed PCFG SET tasks. On average, Transformer outperformed the other two models, but within the two classic neural models, the convolutional model performs better than the LSTM counterpart. In the reported results, two specific architectures, called LSTMS2S and ConvS2S were used. LSTMS2S is a recurrent, bidirectional encoder-decoder model with attention where the encoder and decoder are LSTMs, from the OpenMT-py framework. ConvS2S is a convolution-based sequence-to-sequence model as used in [23]. Several other works [41, 119, 120, 58] have used similar models to conclude that neural sequence models can exploit recursive compositional structure [7, 42] in solving tasks. The related work in compositionality in computer vision indicates similar results. In [53], MLP, CNN, ResNet [32], and relational networks such as WReN [5] and PrediNet [101] are evaluated on substitutivity and productivity tests. Their results indicate that compositional reasoning is challenging for the evaluated models and calls for more sophisticated architectures.
@@ 4.2 Transformer-based Architectures
The compositional capability of large language models is currently a controversial topic. They have been evaluated on general tasks such as arithmetic, logic, and dynamic programming that are compositional by nature. Some of these evaluation efforts conducted in [17] concluded that GPT [84] family Transformers, solve these tasks by reducing them to linearized subgraph matching, without developing true compositional reasoning skills. Moreover, it is shown that, asymptotically, they have architectural limitations in solving highly complex compositional tasks with novel patterns due to error propagation of the composition of erroneous building block functions. There is a substantial gap in the performance of Transformers on in-domain and low-complexity compositional examples versus out-of-domain instances. This indicates they make predictions on shallow reasoning and memorization of similar subgraph patterns seen during training as opposed to reasoning holistically based on true compositional reasoning. The tested tasks were 1) multi-digit multiplication [35], 2) Einstein’s puzzle, which is a constraint satisfaction problem [90], and 3) NP-complete maximum weighted independent set problem [52].
@@
While a series of research papers focused on evaluating the compositional generalization of Transformers [15, 30, 18], some recent research investigated specific architectural factors that can impact the performance of Transformers on compositional tasks, following the claim that Transformers cannot reason compositionally [17]. In [83], five configurations of Transformers were evaluated- 1) type of position encoding, 2) use of copy decoders, 3) model size, 4) weight sharing, and 5) use of intermediate representations for prediction- on several different datasets and benchmarks. The employed tasks were Addition, AdditionNegatives, Reversing, Duplication, Cartesian, Intersection, SCAN-length and SCAN-add-jump, PCFG productivity and systematicity, COGS, and CFQ-mcd1. This work concluded that relative positional encodings usually help, but using embeddings is necessary, and just relative position biases is not enough. Tasks like SCAN and CFQ were not affected by positional embeddings. Tasks like Duplication or PCFG benefit from a copy decoder because it can learn a type of symmetry like learning a certain position of the input. As for model size, it was found that for algorithmic tasks, large models were not helpful. However, for PCFG, large models seemed to outperform their smaller variants. Weight sharing across transformer layers seems to improve accuracy in most tasks. Intermediate representations also improved performance by creating new levels of abstraction that make reasoning easier for solving the end task. Specifically, using intermediate representations achieved state-of-the-art performance on COGS by converting the task from seq2seq to sequence tagging. Using intermediate representation on CFQ, eliminating the need to perform Cartesian products by using triple representations, also showed a significant performance improvement. However, intermediate representations need to be crafted specific to a task and were only tested on COGS and CFQ.
@@
Another type of research deviates from evaluating current conventional architectures and instead focuses on designing a new architecture that can compositionally generalize better.In their model, a multi-modal transformer called GroCoT (Grounded Compositional Transformer) [105] was designed and achieved state-of-the-art results on GSRR and ReaSCAN. The multi-modal transformer from [94] is used as a backbone model with changes to Encoder, Decoder, modified spatial representation, interleaving self-attention, and a modified world state encoding. This work also showed that a single-layer transformer with a single head can ground and compose novel combinations of visual object attributes. They tested the generalizability on RefEx, their proposed evaluation benchmark based on the target identification subtask of ReaSCAN. Another example is adding Pushdown Layers to transformer architecture [78] which was recently presented as a replacement for standard self-attention. The recursive structure of natural language is challenging for self-attention layers to capture due to the lack of an explicit recursive-state tracking mechanism, which Pushdown Layers try to overcome. Pushdown layers have a stack tape that helps them model the recursive state of language. This helps Transformer-based language models softly modulate attention over tokens when predicting new ones. Other architectures derive inspiration from cognitive sciences. In the same line of work, RegularGPT [9] takes inspiration from working memory. It modifies the Transformer architecture to use weight sharing, adaptive depth, and sliding-dilated attention for better length generalization. When tested on the task of natural language extrapolation, it was found that it captures the local windowed attention patterns, which previous work identified as essential for the task. Additionally, it can efficiently model regular languages such as PARITY.
4.3 Neuro-Symbolic Architectures
@@
A rising trend in cutting-edge research on modeling intelligent systems is neuro-symbolic modeling. As the need for general-purpose AI models grows, there is a need for highly compositional models that can reason based on previously trained simpler tasks to do novel and complex ones. One approach in this direction is to use natural language explanations to generate formal specifications that explicitly lay out a compositional task in terms of required simpler steps. The formal specifications then are passed to appropriate engines to solve the problem. A prominent vision understanding model that follows this approach is VisProg [29]. Visprog is a modular neurosymbolic model that can solve various compositional visual reasoning tasks given natural language instruction relying merely on the in-context learning of large language models. It produces modular programs in Python to obtain the solution. This approach provides an interpretable reasoning for how the model derives the solution. These modular programs use built-in modules supported by VisProg such as off-the-shelf neural computer vision models, image preprocessing modules, or Python subroutines, and solve complex tasks without any task-specific training. Another example in this line of work is to generate a formal logical specification of the problem from natural language explanations and pass the logical form to a logical reasoner engine [87]. This work uses large language models such as GPT-3 or GPT-3.5 Turbo, for producing "guides" to solving complex compositional tasks by breaking those down into smaller steps based on a reasoning chain. Similar to this, many recent works focused on different prompting strategies that can be used to solve complex compositional tasks with a modular approach. Examples include Decomposed Prompting [50], which uses a modular approach to decompose a complex task into simpler sub-tasks via prompting and pass on these sub-tasks to LLMs that are capable of solving them. This method allows for the optimization of a prompt for a specific sub-task, which can be further decomposed, or replaced with more effective prompts, trained models, or symbolic functions as necessary. A similar approach for mapping to probabilistic logical reasoning is proposed in [79]. Neuro-symbolic modeling has also been used previously in generative models for concept learning [37] such as in the context of auditory signals for learning evolved combinatorial structure in language.
@@
Another branch of Neuro-symbolic modeling explores leveraging the ability of large language models to do reasoning. This includes casual reasoning with an introduced benchmark, CLadder [46]. Their approach is to provide step-by-step structured prompts as a form of a chain-of-thought strategy called CausalCoT. The chain of thought (COT) conveys the formal symbolic representation of the causal reasoning problem. The CoT prompting is influenced by natural language rationales or reasoning processes [93], which is similar to the use of answer rationales in inducing arithmetic problems [65].
@@
Compositional Neural Architectures. There have been several architectures and theories proposed over the years aimed at modeling compositionality in their design. This is connected to the idea that neural architectures are structurally compositional [62], that is, they take advantage of subroutines to break down complex tasks. One of the earlier examples of this type of architecture includes neural modular networks [2]. Neural modular networks were designed to model the inherent compositionality that exists in linguistic structures. The conceptual modules are built in the neural architecture based on the problem specification. For example, in a visual scene understanding problem, we can place modules for detecting objects, their compositional properties, and relations, which are the main building blocks for abstract reasoning needed for complex scene understanding.
@@
In a similar line of work  [57], a network architecture was built that is compositional in nature and makes it possible to interpret what each part of the network learns. It solves tasks in gSCAN, which has agent navigation tasks in a 2D environment following a natural language command. The neural architecture built in this work, assembles a command-specific network from previously trained modules, modeling the compositional nature of the command (task).
Later research showed, that in Neural Module Networks, it is hard to make the designed modules faithful to expressing the concepts that they are designed for, despite the overall network achieving high accuracy for the target task [112].
@@
Neurocompositional Computing. Neuro-symbolic modeling has been motivated by its connection to neurocompositional computing. The term "neurocompositional computing" was coined in [109]. It defines a type of computing that underlies human cognition as argued in contemporary cognitive science theories in [108] and incorporates principles of Continuity and Compositionality. The Continuity Principle states that the encoding and processing of information should be continuous, that is, represented by real numbers that vary continuously and can be changed by arbitrarily small amounts. The Compositionality Principle states that larger, more complex structures can be decoded on the basis of smaller, simpler, and familiar building blocks. According to the Central Paradox of Cognition, the human brain follows both a continuous neural computing structure and a discrete compositional-structure computer. Following this theory, neuro-symbolic models that are both continuous and discrete in architecture seem like the ideal approach to modeling compositional behavior in computing architectures [110].
5 Theoretical Findings
We overview the theoretical results that investigate the compositional generalization of classical neural networks, transformers, and modern language models.

### v2

@@ 3.2 Productivity or Length Generalization
PCFG SET. For the Productivity test (Hupkes et al., 2020), the data is split based on sequence lengths. The model is tested on sequences longer than the ones observed during training. For example, in a grammar context, if the model has been trained on the syntax entity-relation-entity, it will be tested on a longer, nested version of this concept, entity-relation-(entity-relation-entity).
CFQ. This is a dataset of Compositional Freebase Questions (CFQ). It is a natural language question-answering task  (Keysers et al., 2020), focusing on semantic parsing, with a SPARQL query against the Freebase knowledge base. Questions are generated at varying levels of complexity. Various splits are available based on input length, output length, input pattern, or output pattern. All these splits aim to maximize compound divergence while minimizing atom divergence. This task appears to test both systematicity and productivity to some extent, although not explicitly. However, it cannot identify specific areas where a model’s compositional behavior may be deficient.
COGS. While some splits of this dataset are explained in Section 3.1, it also contains a split for length generalization. Category 3 (i.e. Deeper Recursion), one of the splits of this dataset,  (Kim & Linzen, 2020) is a test for length generalization by increasing the length of the input sequence recursively during testing. Input sequences are generated using nesting of phrases that are longer than those seen during training.
@@ 3.3 Other Generalization Criteria
To the best of our knowledge, PCFG SET (Hupkes et al., 2020) is the only benchmark that evaluates the other three additional criteria. The Substitutivity or synonymity test uses an input sequence with an atomic unit replaced by a synonymous atomic unit to evaluate how the model prediction changes. Localism is tested by using input sequences composed of smaller sequences A and B. The model is used to translate the full sequences first and then forced to process A and B separately. The outputs of these two experiments are compared to evaluate how local versus global the model is in its compositional reasoning. Overgeneralization test evaluates the model’s results on input sequences that do not conform to the general rules of the dataset, that is, input sequences that are exceptions to the dataset rules. For example, the acquisition of past-tense forms such as the common “ed” ending (open-opened) versus more uncommon forms such as break to broke.
Name Text MM. Sys. Prod. Subst. Loc. Overgen. References PCFG SET  Csordás et al. (2021) CLUTRR  Gontier et al. (2020),  Minervini et al. (2020)  Sinha et al. (2019) SQOOP  D’Amario et al. (2021), Bahdanau et al. (2019) CFQ  Furrer et al. (2021)Herzig et al. (2021),  Liu et al. (2021)Cao et al. (2022) SCAN  Korrel et al. (2019),  Nye et al. (2020),  Dessì & Baroni (2019) COGS  Haurilet et al. (2019)Wu et al. (2023),  Klinger et al. (2024) gSCAN  Gao et al. (2020),  Spilsbury et al. (2024) ReaSCAN  Kamali & Kordjamshidi (2023) CREPE  Lin et al. (2024)Singh et al. (2023) KiloGram  Kojima et al. (2023)Gui et al. (2023) Ji et al. (2022) CLEVR  Bahdanau et al. (2020), Johnson et al. (2016),  Niemeyer & Geiger (2021) MIT-States  Xu et al. (2023a), Naeem et al. (2021)
@@
4 Empirical Findings: Compositional Learning Models
Traditional symbolic AI models naturally support compositional reasoning using classical logic applied to formal language (Szabó, 2022), formal verification (Giannakopoulou et al., 2018), and more. First-order logic can express objects, their properties, and complex compositional relations. Logical operations like conjunction, disjunction, and implication can express compositional structures on which inference rules are applied, supporting complex compositional reasoning (Porto, 2002). Another classical symbolic formalism includes grammars (Chomsky, 1965), which can express and generate complex compositional structures. However, dealing with noisy and uncertain data is hard with pure symbolic AI. Moreover, probabilistic augmentations and structured output prediction models have been able to explicitly model structural dependencies and support compositional reasoning based on their learned complex patterns from the data (Pearl, 1988). Nevertheless, scalability becomes a challenge for training and inference as the structural dependencies and the number of correlated variables increase. Given these long-lasting challenges of traditional models of compositionality, current neural models have shown success in both scalability and dealing with noisy and sensory data (OpenAI, 2024). Especially in modern large language models, complex compositional patterns can be memorized and resemble compositional reasoning. In the rest of this section, we overview the research focused on the development, design, and empirical evaluation of different types of neural models for compositional reasoning. We relate the type of empirical evaluations to the cognitive aspects of compositionality that they are testing. While some of these models utilize tasks and datasets covered in Section 3, others have their own tasks, specific to the problem of their choice.
@@ 4.1 Basic Neural Models
In Hupkes et al. (2020), different neural models are tested on a set of compositional learning tasks. They evaluate Long short-term memory (LSTM) Networks (Hochreiter & Schmidhuber, 1997), Convolutional Neural Networks (CNN) (LeCun et al., 1989), and Transformers for sequence-to-sequence language processing tasks on their proposed PCFG SET tasks including systematicity, productivity, substitutivity, localism and overgeneralization. On average, Transformer outperformed the other two models, but within the two classic neural models, the convolutional model performs better than the LSTM counterpart. In the reported results, two specific architectures, called LSTMS2S and ConvS2S were used. LSTMS2S is a recurrent, bidirectional encoder-decoder model with attention where the encoder and decoder are LSTMs, from the OpenMT-py framework. ConvS2S is a convolution-based sequence-to-sequence model as used in Gehring et al. (2017). Several other works (Hupkes et al., 2018; Zheng & Lapata, 2021; 2022; Lake & Baroni, 2017) have used similar models to conclude that neural sequence models can exploit recursive compositional structure (Bowman et al., 2015; Irsoy & Cardie, 2014) in solving tasks. The related work in compositionality in computer vision indicates similar results. In Klinger et al. (2020), MLP, CNN, ResNet (He et al., 2015), and relational networks such as WReN (Barrett et al., 2018) and PrediNet (Shanahan et al., 2020) are evaluated on PCFG SET substitutivity and productivity tests. Their results indicate that compositional reasoning is challenging for the evaluated models and calls for more sophisticated architectures.
@@ 4.2 Transformer-based Architectures
The compositional capability of large language models is currently a controversial topic. They have been evaluated on general tasks such as arithmetic, logic, and dynamic programming that are compositional by nature Nafar et al. (2024a; b). Some of these evaluation efforts conducted in Dziri et al. (2023) concluded that GPT (OpenAI, 2024) family models, solve these tasks by reducing them to linearized subgraph matching, without developing true compositional reasoning skills. Moreover, it is shown that, asymptotically, they have architectural limitations in solving highly complex compositional tasks with novel patterns due to error propagation of the composition of erroneous building block functions. There is a substantial gap in the performance of Transformers on in-domain and low-complexity compositional examples versus out-of-domain instances. The tested tasks were 1) multi-digit multiplication (Hiebert, 2013), 2) Einstein’s puzzle, which is a constraint satisfaction problem (Prosser, 1993), and 3) NP-complete maximum weighted independent set problem (Kleinberg & Tardos, 2005). These tasks are mostly aimed at testing systematicity and productivity. Their results indicate that transformers make predictions on shallow reasoning and memorization of similar subgraph patterns seen during training as opposed to reasoning holistically based on true compositional reasoning. This hypothesis also aligns with the findings presented in Chang & Bisk (2024). They conduct experiments with counting problems, a basic form of length generalization tasks. While transformers can count in observed cases, they dramatically fail to perform out-of-domain for the same task, indicating that transformers rely on memorizing observations. The results on transformers often extend to transformer-based language models. As shown in this work (Anil et al., 2022), fine-tuned transformer-based large language models lack generalization capabilities irrespective of the model size, when tested on other length generalization tasks such as parity and boolean variable assignment. This is due to the transformer’s tendency to learn non-sequential patterns that do not apply to longer sequences. However, combining a pretrained LLM’s in-context learning ability with scratchpad prompting significantly improves performance on longer sequences. This also implies that despite having access to an infinite data pool, LLMs can potentially learn some tasks better through in-context learning than finetuning, confirmed by theoretical work on LLMs explained in Section 5.
@@
While a series of research papers focused on evaluating the compositional generalization of Transformers (Dehghani et al., 2019; Hahn, 2020; Feng et al., 2023) and complex reasoning Mirzaee et al. (2021); Mirzaee & Kordjamshidi (2022), some recent research investigated specific architectural factors that can impact the performance of Transformers on compositional tasks, following the claim that Transformers cannot reason compositionally (Dziri et al., 2023). In Ontanon et al. (2022), five configurations were evaluated on several different datasets and benchmarks, by varying five properties of Transformers including, 1) type of positional encoding, 2) use of copy decoders, 3) model size, 4) weight sharing, and 5) use of intermediate representations for prediction. The employed tasks were Addition, AdditionNegatives, Reversing, Duplication, Cartesian, Intersection, SCAN-length and SCAN-add-jump, PCFG productivity and systematicity, COGS, and CFQ-mcd1. This work concluded that relative positional encodings usually help, but using embeddings is necessary, and merely relative position biases are not sufficient. Tasks like SCAN and CFQ were not affected by positional embeddings. Tasks like Duplication or PCFG benefit from a copy decoder because it can learn a type of symmetry like learning a certain position of the input. As for model size, it was found that for algorithmic tasks, large models did not make a difference. However, for PCFG, large models seemed to outperform their smaller variants. Weight sharing across transformer layers seems to improve accuracy in most tasks. Intermediate representations also improved performance by creating new levels of abstraction that make reasoning easier for solving the end task. Specifically, using intermediate representations achieved state-of-the-art performance on COGS by converting the task from seq2seq to sequence tagging. Using intermediate representation on CFQ, eliminating the need to perform Cartesian products by using triple representations, also showed a significant performance improvement. However, intermediate representations need to be crafted specifically for a task and were only tested on COGS and CFQ. Other works (Ruoss et al., 2023; Kazemnejad et al., 2023; Chang & Bisk, 2024) also investigate the effect of different positional encodings on the performance of Transformers for specifically length generalization. Kazemnejad et al. (2023) concludes that commonly used positional encodings such as ALiBi and APE are not well suited for downstream tasks but transformers without any positional encoding (NoPE) outperform other explicit positional encoding methods when testing decoder-only transformers on tasks such as SCAN and copying and reversing. Despite this, Chang & Bisk (2024) presents compelling evidence suggesting that each type of positional encoding demonstrates robustness on unique length generalization tasks. Consequently, the performance of Transformers on compositional tasks varies depending on the selected positional encoding. The authors even suggest integrating multiple positional encodings into Transformers for better compositionality by combining different strengths of those encodings.
@@
Another type of research deviates from evaluating current conventional architectures and instead focuses on designing a new architecture that can compositionally generalize better. One of such models, a multi-modal transformer called GroCoT (Grounded Compositional Transformer) (Sikarwar et al., 2022), was designed and achieved state-of-the-art results on GSRR and ReaSCAN. Another multi-modal transformer from Qiu et al. (2021) is used as a backbone model with changes to Encoder, Decoder, modified spatial representation, interleaving self-attention, and a modified world state encoding. This work also showed that a single-layer transformer with a single head can ground and compose novel combinations of visual object attributes. They tested the generalizability on RefEx, that is on grounding referring expressions and, their proposed evaluation benchmark based on the target identification subtask of ReaSCAN. Another customized architecture example is adding Pushdown Layers to transformer architecture (Murty et al., 2023) which was recently presented as a replacement for standard self-attention. The recursive structure of natural language is challenging for self-attention layers to capture due to the lack of an explicit recursive-state tracking mechanism, which Pushdown Layers try to overcome. Pushdown layers have a stack tape that helps them model the recursive state of language. This helps Transformer-based language models softly modulate attention over tokens when predicting new ones. Other architectures derive inspiration from cognitive sciences. In the same line of work, RegularGPT (Chi et al., 2023) takes inspiration from working memory. It modifies the Transformer architecture to use weight sharing, adaptive depth, and sliding-dilated attention for better length generalization. When tested on the task of natural language extrapolation, it was found that it captures the local windowed attention patterns, which previous work identified as essential for the task. Additionally, it can efficiently model regular languages such as PARITY.
4.3 Neurosymbolic Architectures
@@
A rising trend in cutting-edge research on modeling intelligent systems is neurosymbolic modeling. As the need for general-purpose AI models grows, there is a need for highly compositional models that can reason based on previously trained simpler tasks to do novel and complex ones. Although not explicitly mentioned in this research, they mainly address systematicity and productivity. One approach in this direction is to use natural language explanations to generate formal specifications that explicitly lay out a compositional task in terms of required simpler steps. The formal specifications then are passed to appropriate engines to solve the problem. A prominent vision understanding model that follows this approach is VisProg (Gupta & Kembhavi, 2023). VisProg is a modular neurosymbolic model that can solve various compositional visual reasoning tasks given natural language instruction relying merely on the in-context learning of large language models. It produces modular programs in Python to obtain the solution. This approach provides an interpretable reasoning for how the model derives the solution. These modular programs use built-in modules supported by VisProg such as off-the-shelf neural computer vision models, image preprocessing modules, or Python subroutines, and solve complex tasks without any task-specific training. Another example in this line of work is to generate a formal logical specification of the problem from natural language explanations and pass the logical form to a logical reasoner engine (Poesia et al., 2023). This work uses large language models such as GPT-3 or GPT-3.5 Turbo, for producing “guides” to solving complex compositional tasks by breaking those down into smaller steps based on a reasoning chain. Similar to this, many recent works focused on different prompting strategies that can be used to solve complex compositional tasks with a modular approach. Examples include Decomposed Prompting (Khot et al., 2023), which uses a modular approach to decompose a complex task into simpler sub-tasks via prompting and pass on these sub-tasks to LLMs that are capable of solving them. This method allows for the optimization of a prompt for a specific sub-task, which can be further decomposed, or replaced with more effective prompts, trained models, or symbolic functions as necessary. A similar approach for mapping to probabilistic logical reasoning is proposed in Nafar et al. (2024b). Neurosymbolic modeling has also been used previously in generative models for concept learning (Hofer et al., 2021) such as in the context of auditory signals for learning evolved combinatorial structure in language.
@@
Another branch of Neurosymbolic modeling explores leveraging the ability of large language models to do reasoning. This includes casual reasoning with an introduced benchmark, CLadder (Jin* et al., 2023). Their approach is to provide step-by-step structured prompts as a form of a chain-of-thought strategy called CausalCoT. The chain of thought (COT) conveys the formal symbolic representation of the causal reasoning problem. The CoT prompting is influenced by natural language rationales or reasoning processes (Qiao et al., 2023), which is similar to the use of answer rationales in inducing arithmetic problems (Ling et al., 2017).
@@
Compositional Neural Architectures. Over the years, several architectures and theories have been proposed to model compositionality in their design. This aligns with the idea that neural architectures are structurally compositional (Lepori et al., 2023), meaning they leverage subroutines to break down complex tasks. One of the earlier examples of this type of architecture includes neural modular networks (Andreas et al., 2016). Neural modular networks were designed to model the inherent compositionality that exists in linguistic structures. The conceptual modules are built in the neural architecture based on the problem specification. For example, in a visual scene understanding problem, we can place modules for detecting objects, their compositional properties, and relations, which are the main building blocks for abstract reasoning needed for complex scene understanding.
@@
In a similar line of work (Kuo et al., 2020), a network architecture was built that is compositional in nature and makes it possible to interpret what each part of the network learns. It solves tasks in gSCAN, which has agent navigation tasks in a 2D environment following a natural language command. The neural architecture built in this work assembles a command-specific network from previously trained modules, modeling the compositional nature of the command (task). Later research showed that in Neural Module Networks, it is hard to make the designed modules faithful to expressing the concepts that they are designed for. This is despite the overall network achieving high accuracy for the target task (Subramanian et al., 2020).
Given the importance of compositional learning in lifelong and continual learning, neurosymbolic approaches have been a suitable framework to address continuality through combining program synthesis and neural modeling. One such example is HOUDINI (Valkov et al., 2018), which is a neurosymbolic framework for lifelong learning of tasks combining perception and procedural reasoning such as counting, summing, and shortest-path computation. They use program synthesis to search over networks described as typed functional programs for the given task, whose parameters are then tuned end-to-end based on stochastic gradient descent. Another example is the Logic-Enhanced Foundation Model (LEFT) (Hsu et al., 2023), a framework designed to learn grounding and reasoning across domains using a differentiable, domain-independent, first-order logic-based program executor. It addresses the lack of generalizability across domains seen in related work, such as VisProg, which we previously discussed, including limitations in generalizing concepts from 2D to 3D images.
@@
Neurocompositional Computing. Neurosymbolic modeling has been motivated by its connection to neurocompositional computing. The term “neurocompositional computing” was coined in Smolensky et al. (2022b). It defines a type of computing that underlies human cognition as argued in contemporary cognitive science theories in Smolensky & Legendre (2006) and incorporates principles of Continuity and Compositionality. The Continuity Principle states that the encoding and processing of information should be continuous, that is, represented by real numbers that vary continuously and can be changed by arbitrarily small amounts. The Compositionality Principle states that larger, more complex structures can be decoded on the basis of smaller, simpler, and familiar building blocks. According to the Central Paradox of Cognition, the human brain follows both a continuous neural computing structure and a discrete compositional-structure computer. Following this theory, neurosymbolic models that are both continuous and discrete in architecture seem like the ideal approach to modeling compositional behavior in computational models (Smolensky et al., 2022a).
5 Theoretical Findings: Mathematical Formulations of Compositionality
Theoretical analysis is fundamental for deepening our understanding of the compositionality of learning models. It can reveal intriguing and previously uncovered information that experimental analysis may overlook. Many research works have proposed diverse approaches for investigating the compositionality of learning models. We highlight three different approaches, including a mathematical framework for defining compositionality (Ram et al., 2023), exploring the upper-bounds of expressivity that relate to compositionality (Merrill & Sabharwal, 2023), and analyzing error-bounds to demonstrate the model’s limitations in solving compositional learning problems (Dziri et al., 2023). In the rest of this section, we provide a detailed overview of these cases and explain the theoretical results on compositional generalization of classical neural networks, transformers, and modern language models. We also relate the mentioned techniques to aspects of compositionality when applicable.
5.1 Classical Neural Network